d4 11,000 -4,00015 O 10,700 O 5,400 O 3,800 3,500 O 14,200

Question 5 5 pts Use the table below to determine the expected value of perfect information. The probabilities for the states of nature are s1-0.20, s2-0.30 and s3-0.50 State of Nature s2 10,000 -5,000 15,000 15,000 9,000 10,000 20,0005,000 10,000 11,0004,000 15,000 s1 S3 Decision d1 d2 d3 d4 10700
d4 11,000 -4,00015 O 10,700 O 5,400 O 3,800 3,500 O 14,200
Answer #1

3500

27755 078 1 0 1 59 54 0 2 1 1234

EVPI= 0.2*20 + 0.3* 9+ 0.5*15 - 10.7= 14.2-10.7= 3.5 (3500)